OpenStudy (anonymous):

(14xy)^-8/9 write this expression with positive exponents

OpenStudy (anonymous):

Send that to denominator

OpenStudy (anonymous):

\[a^{-4} =1/a^4\]

OpenStudy (anonymous):

\[\frac{1}{(14xy)^{\frac{8}{9}}}\]